OpenStudy (anonymous):

Find f '(a). f(t) = (9t + 7)/(t + 2)

OpenStudy (turingtest):

f'(t)=[9(t+2)-1(9t+7)]/[(t+2)^2]=-5/[(t+2)^2] f'(a)=-5/(a^2+4a+4)